# The Cohen–Lenstra moments over function fields via the stable homology of non-splitting Hurwitz spaces

### v2

@@ 2.3. Covers split completely at infinity
First, we describe the case that $n$ is even. Any hyperelliptic curve in characteristic not $2$ which is unramified over $\infty$ can be uniquely written in the form $f(x)=\alpha x^{n}+a_{n-1}x^{n-1}+\cdots+a_{1}x+a_{0}$ where either $\alpha=1$ or $\alpha\in\mathbb{F}_{q}^{\times}$ is a fixed non-square. Such a curve is split completely over infinity if and only if its fiber over $\infty\in\mathbb{P}^{1}_{\mathbb{F}_{q}}(\mathbb{F}_{q})$ consists of two $\operatorname{Spec}\mathbb{F}_{q}$ points. The fiber over $\infty$ of this curve can be identified with the fiber over 00 of the curve $z^{2}=\alpha+a_{n-1}u+\cdots+a_{1}u^{n-1}+a_{0}u^{n}$, where we have set $u=1/x$ and then $z=y/x^{n/2}$. This is the same as $\operatorname{Spec}\mathbb{F}_{q}[y]/(y^{2}-\alpha)$, which will be two copies of $\mathbb{F}_{q}$ when $\alpha$ is a square and a single copy of $\mathbb{F}_{q}^{2}$ when $\alpha$ is a non-square. Therefore, the curves where $\alpha$ is a square (which are all isomorphic to such curves with $\alpha=1$) precisely correspond to those completely split over $\infty$.
@@
We next consider the case that $n$ is odd. As above, we can still write our hyperelliptic curve in the form $f(x)=\alpha x^{n}+a_{n-1}x^{n-1}+\cdots+a_{1}x+a_{0}$ with $\alpha$ either $1$ or a fixed non-square. Changing coordinates to $u=1/x$, we can write the curve in the form $z^{2}=u(\alpha+a_{n-1}u+\cdots+a_{1}u^{n-1}+a_{0}u^{n})$ where $u=1/x$ and $z=y/x^{n+1/2}$. Let $X$ denote the proper regular curve corresponding to the above equation. We can identify the stack quotient $[X/(\mathbb{Z}/2\mathbb{Z})]$ Zariski locally around the residual gerbe at $\infty$ with a Zariski open of the root stack $\mathscr{P}$, which is a root stack of $\mathbb{P}^{1}$ over the divisor $\infty$, so $\mathscr{P}$ has a single stacky point of order $2$.
@@
Let ~$\widetilde{\infty}:\operatorname{Spec}\mathbb{F}_{q}\to\mathscr{P}$ denote the map which factors through the residual gerbe $B\mu_{2}$ at infinity and corresponds to the map $\operatorname{Spec}\mathbb{F}_{q}\to B\mu_{2}$ associated to the trivial double cover $\operatorname{Spec}\mathbb{F}_{q}\coprod\operatorname{Spec}\mathbb{F}_{q}\to\operatorname{Spec}\mathbb{F}_{q}$. We wish to show that the fiber product $X\times_{\mathscr{P},\infty}\operatorname{Spec}\mathbb{F}_{q}$ consists of two copies of $\operatorname{Spec}\mathbb{F}_{q}$. The fiber product above is identified with the fiber product $X\times_{\mathscr{P}}\mathbb{P}^{1}\times_{\mathbb{P}^{1},\infty}\operatorname{Spec}\mathbb{F}_{q}$, where the map $\mathbb{P}^{1}\to\mathscr{P}$ is the double cover branched over 00. (In particular, on coarse spaces $\mathbb{P}^{1}\to\mathscr{P}$ corresponds to the map given in local coordinates by $x\mapsto x^{2}$.) The above fiber product $X\times_{\mathscr{P}}\mathbb{P}^{1}$ is a normal curve in a neighborhood over $\infty$, and hence it is, locally around $\infty$, the normalization of $X\times_{\mathbb{P}^{1}_{u}}\mathbb{P}^{1}_{t}$, where again the map $\mathbb{P}^{1}_{t}\to\mathbb{P}^{1}_{u}$ is induced by the double cover $u\mapsto t^{2}$. In a neighborhood of $\infty$, we can express the projection map $X\times_{\mathbb{P}^{1}_{u}}\mathbb{P}^{1}_{t}\to X$ as
@@ 3.1.6. Chains functor
We have a symmetric monoidal functor $\tilde{C}_{*}(-;\mathbb{Q}):\operatorname{Spc}_{+}\to\operatorname{Sp}$, taking $A$ to its reduced chains with coefficients in $\mathbb{Q}$, with the comultiplication coming from the diagonal map, which is dual to the cup product. Composing $\tilde{C}_{*}(-;\mathbb{Q})$ with the symmetric monoidal functor $(-)_{+}$, we get the symmetric monoidal functor $C_{*}(-;\mathbb{Q}):\operatorname{Spc}\to\operatorname{Sp}$. For $X$ a space, $H_{i}(X;\mathbb{Q})=\pi_{i}C_{*}(X;\mathbb{Q})$.
@@ 3.3. Localization and completion
[…]
@@
There are natural $\mathbb{E}_{1}$-ring maps $R\to R^{\wedge}_{r}$, $R\to R[r^{-1}]$ induced by taking the map induced on endomorphism rings of the $R$-module $R$ via the completion and localization functor. Note also that $R[r^{-1}]=R[(r^{n})^{-1}]$ and $R^{\wedge}_{r}=R^{\wedge}_{r^{n}}$ for each $n\geq 1$.
@@
Because $x^{m}$ maps to 00 in the component $R[(1-e)^{-1}]\simeq R^{\wedge}_{x}$, the kernel of surjective map $\pi_{*}R\to\pi_{*}R^{\wedge}_{x}$ contains two sided ideal generated by $x^{m}$. From the above, this kernel is equal to $\pi_{*}R[x^{-1}]$. Since $\pi_{*}R[x^{-1}]$ is contained in the two sided ideal generated by $x^{m}$ since any $y\in\pi_{*}R[x^{-1}]$ can be written as $yx^{-m}x^{m}$, we obtain that the kernel is equal to the two sided ideal generated by $x^{m}$. ∎
@@ 3.4. Tools for homological epimorphisms
Using § 3.4, we learn that $S\to S^{\prime}$ is a homological epimorphism. Letting $I$ be the fiber of $S\to S^{\prime}$, we have $I\otimes_{S}S^{\prime}=0$. $I$ is the fiber of a map of connective $\mathbb{E}_{1}$-rings, so it is $-1$-connective. We will inductively show on $i\geq-1$ that $\pi_{i}I$ is 00. The inductive hypothesis implies that $I$ is $i$-connective, so and since $S\to S^{\prime}$ is an equivalence on $\pi_{0}$, we claim
@@ 3.5. Nilpotence of actions
[…]
@@
In particular, if a natural endomorphism $g$ of the identity functor acts by 00 on $X$ and $Y$, then $g^{2}$ acts by 00 on $\operatorname{cof}(f)$ for any map $X\to Y$.
@@
For the first claim, the composite $\operatorname{cof}(f)\to\operatorname{cof}(f^{\prime})\to\Sigma X^{\prime}$ is null since $X\to X^{\prime}$ is null, so we obtain a factorization of $\operatorname{cof}(f)\to\operatorname{cof}(f^{\prime})$ through $Y^{\prime}$. Composing with the second horizontal map, since $Y^{\prime}\to Y^{\prime\prime}$ is 00, we obtain a nullhomotopy of the desired map.
For the second claim, we apply the first claim when $X=X^{\prime}=X^{\prime\prime},Y=Y^{\prime}=Y^{\prime\prime}$ and $f=f^{\prime}=f^{\prime\prime}=g$. ∎
@@
Let ${\mathcal{C}}$ be an exactly monoidal stable $\infty$-category, let $f:\operatorname{\mathbb{1}}\to\operatorname{\mathbb{1}}$ be a map in ${\mathcal{C}}$, where $\operatorname{\mathbb{1}}$ is the monoidal unit. Then the map $\operatorname{cof}(f)\otimes f^{\otimes 2}$ is null.
In particular, if $R$ is an $\mathbb{E}_{1}$-algebra, $f:R\to\Sigma^{i}R$ a bimodule map, and $M\in\operatorname{Mod}_{\operatorname{Sp}}(R)$, then $f^{2}$ acts by 00 on $\operatorname{cof}f\otimes_{R}M$ as a map of $R$-modules.
@@
[…]
@@ 4.1. Background on racks
We can identify the connected components of the pointed Hurwitz space $\operatorname{Hur}^{c}_{n}$ with orbits of the $B_{n}$ action on $c^{n}$. Under this identification, for $x_{1},\ldots,x_{n}\in c$ we use $[x_{1}]\cdots[x_{n}]$ to denote the connected component of $\operatorname{Hur}^{G,c}_{n,\mathbb{C}}$ corresponding to the $B_{n}$ orbit of the tuple $(x_{1},\ldots,x_{n})$.
@@ 4.2. The main stability result
First, note that there is an identification between the pointed Hurwitz space $\operatorname{Hur}_{n}^{G,c}$ and $\operatorname{Hur}_{n}^{c}$ as both can be described as the homotopy quotient $(c^{n})_{hB_{n}}$, see § 1.4.2. Let us fix a value of $i\geq 0$. By [EVW16, Theorem 6.1], there exist constants $I,J$, and $D$ [^4] In fact, one may take $D=1$ by [EL23, Proposition A.3.1]. so that for any $i\geq 0$ and $n>Ii+J$ and $U:=\sum_{h\in c}\alpha_{h}^{D\operatorname{ord}(h)}$, $\pi_{i}A_{n}$ agrees with $\pi_{i}A[U^{-1}]_{n}$.
@@ 4.3. Centrality
Informally, in the case of Hurwitz spaces associated to groups, the reason for this stronger form of centrality is that there is a homotopy $[\alpha][\beta]\simeq[\beta][\beta^{-1}\alpha\beta]$ obtained by changing the direction from which one concatenates $G$-covers. Thus if $\beta$ centralizes $c$, this identifies left and right multiplication by $\beta$. The above argument is not enough to make multiplication by $\beta$ into a bimodule map, but it can be made as such. We now argue this below and further generalize from groups to racks:
@@
Let $m$ be such that $\gamma\in\operatorname{Hur}_{m}^{c}$. Consider $m_{!}(*)$, where $m$ is considered as the inclusion of the base point $*\to BB_{m}$. Explicitly, $m_{!}(*)$ is the local system which is $B_{m}$ in degree $m$ and empty otherwise. There is a left module map $(c^{i})_{i\in\mathbb{N}}\otimes m_{!}(*)\to(c^{i})_{i\in\mathbb{N}}$, given by right multiplying by the class $\gamma$ on the identity element of $B_{m}$. Here we realize $(c^{i})_{i\in\mathbb{N}}\otimes m_{!}(*)$ as a bimodule via the $\mathbb{E}_{2}$-monoidal structure, i.e using the twist map $(c^{i})_{i\in\mathbb{N}}\otimes m_{!}(*)\to m_{!}(*)\otimes(c^{i})_{i\in\mathbb{N}}$. The fact that this multiplication map is a bimodule map follows if the diagram
@@
Let $\gamma\in\pi_{0}\operatorname{Hur}^{c}$ be central. Then the associated class $[\gamma]\in\pi_{0}A$ is $\mathbb{E}_{2}$-central.
@@ 4.4. Decomposition of the comparison map
[…]
@@
With notation as in § 4.4, let $c$ be a finite nonempty non-splitting rack and let $c^{\prime}\subset c$ be a subrack. In $A[\alpha_{c^{\prime}}^{-1}]$, any element $\alpha_{\beta}$ with $\beta\in c-c^{\prime}$ satisfies the conditions of § 3.3. Thus for an $A$-module $M$, we can form $M[\alpha_{c^{\prime}}^{-1}]^{\wedge}_{\alpha_{c-c^{\prime}}}$ as in § 3.3.
@@
is an equivalence, where $\coprod_{i=1}^{|c-c^{\prime}|}\beta_{i}=c-c^{\prime}$. Here, we have used § 3.4 to commute the localization with the quotienting, and § 4.3 to make sense of the iterated quotient. Namely, because $\alpha_{\beta_{i}}^{\operatorname{ord}(\beta_{i})}$ is a bimodule map by § 4.3, we can tensor with the $A-A$-bimodule $(A/\alpha^{\operatorname{ord}(\beta_{1})}_{\beta_{1}})$ to make sense of the tensor product above as an $A$-module.
Note that on the source and target of this map, for each $\beta\in c-c^{\prime}$, $\alpha_{\beta}^{\operatorname{ord}(\beta)n_{\beta}}$ acts by 0 for some $n_{\beta}$. This follows from applying § 3.5 to see that after quotienting out by $\alpha_{\beta}^{\operatorname{ord}(\beta)}$, $\alpha_{\beta}^{2\operatorname{ord}(\beta)}$ acts by 00, and using § 3.5 to see that after each further quotient, the smallest power of $\alpha_{\beta}^{\operatorname{ord}(\beta)}$ that acts by 00 gets multiplied by at most $2$.
If $c^{\prime}$ is empty, then we claim that both the source and target of the map (4.1) are 00. We claim that some power of $U$ acts by 00 on the source and target, since a large enough power of $U$, $U^{j}$, is in the two sided ideal of $\pi_{0}A$ generated by $\alpha_{\beta}^{\operatorname{ord}(\beta)n_{\beta}}$. This is because by the pigeonhole principle, one can see that any element of $\pi_{0}\operatorname{Hur}^{G,c}_{n}$, written as a product of elements in $c$ for $n$ large enough contains $\max_{\beta\in c}(\operatorname{ord}(\beta)n_{\beta})$ many copies of a single $\beta\in c$. By pulling these to the left using the braid group action, we see the element is divisible by $\alpha_{\beta}^{\operatorname{ord}_{\beta}n_{\beta}}$. Since $U^{j}$ is in the ideal generated by the smallest powers of $\alpha_{\beta}^{\operatorname{ord}(\beta)}$ that act trivially, and hence it acts trivially. Since $U$ acts both as an isomorphism and nilpotently, the source and target of (4.1) with $c^{\prime}=\emptyset$ are both 00.

### v3

@@ 2.3. Covers split completely at infinity
First, we describe the case that $n$ is even. Any hyperelliptic curve in characteristic not $2$ which is unramified over $\infty$ can be uniquely written in the form $f(x)=\alpha x^{n}+a_{n-1}x^{n-1}+\cdots+a_{1}x+a_{0}$ where either $\alpha=1$ or $\alpha\in\mathbb{F}_{q}^{\times}$ is a fixed non-square. Such a curve is split completely over infinity if and only if its fiber over $\infty\in\mathbb{P}^{1}_{\mathbb{F}_{q}}(\mathbb{F}_{q})$ consists of two $\operatorname{Spec}\mathbb{F}_{q}$ points. The fiber over $\infty$ of this curve can be identified with the fiber over 0 of the curve $z^{2}=\alpha+a_{n-1}u+\cdots+a_{1}u^{n-1}+a_{0}u^{n}$, where we have set $u=1/x$ and then $z=y/x^{n/2}$. This is the same as $\operatorname{Spec}\mathbb{F}_{q}[y]/(y^{2}-\alpha)$, which will be two copies of $\mathbb{F}_{q}$ when $\alpha$ is a square and a single copy of $\mathbb{F}_{q}^{2}$ when $\alpha$ is a non-square. Therefore, the curves where $\alpha$ is a square (which are all isomorphic to such curves with $\alpha=1$) precisely correspond to those completely split over $\infty$.
@@
We next consider the case that $n$ is odd. As above, we can still write our hyperelliptic curve in the form $f(x)=\alpha x^{n}+a_{n-1}x^{n-1}+\cdots+a_{1}x+a_{0}$ with $\alpha$ either $1$ or a fixed non-square. Changing coordinates to $u=1/x$, we can write the curve in the form $z^{2}=u(\alpha+a_{n-1}u+\cdots+a_{1}u^{n-1}+a_{0}u^{n})$ where $u=1/x$ and $z=y/x^{n+1/2}$. Let $X$ denote the proper regular curve corresponding to the above equation. We can identify the stack quotient $X/(\mathbb{Z}/2\mathbb{Z})]$ Zariski locally around the residual gerbe at $\infty$ with a Zariski open of the root stack $\mathscr{P}$, which is a root stack of $\mathbb{P}^{1}$ over the divisor $\infty$, so $\mathscr{P}$ has a single stacky point of order $2$.
@@
Let $\widetilde{\infty}:\operatorname{Spec}\mathbb{F}_{q}\to\mathscr{P}$ denote the map which factors through the residual gerbe $B\mu_{2}$ at infinity and corresponds to the map $\operatorname{Spec}\mathbb{F}_{q}\to B\mu_{2}$ associated to the trivial double cover $\operatorname{Spec}\mathbb{F}_{q}\coprod\operatorname{Spec}\mathbb{F}_{q}\to\operatorname{Spec}\mathbb{F}_{q}$. We wish to show that the fiber product $X\times_{\mathscr{P},\infty}\operatorname{Spec}\mathbb{F}_{q}$ consists of two copies of $\operatorname{Spec}\mathbb{F}_{q}$. The fiber product above is identified with the fiber product $X\times_{\mathscr{P}}\mathbb{P}^{1}\times_{\mathbb{P}^{1},\infty}\operatorname{Spec}\mathbb{F}_{q}$, where the map $\mathbb{P}^{1}\to\mathscr{P}$ is the double cover branched over 0. (In particular, on coarse spaces $\mathbb{P}^{1}\to\mathscr{P}$ corresponds to the map given in local coordinates by $x\mapsto x^{2}$.) The above fiber product $X\times_{\mathscr{P}}\mathbb{P}^{1}$ is a normal curve in a neighborhood over $\infty$, and hence it is, locally around $\infty$, the normalization of $X\times_{\mathbb{P}^{1}_{u}}\mathbb{P}^{1}_{t}$, where again the map $\mathbb{P}^{1}_{t}\to\mathbb{P}^{1}_{u}$ is induced by the double cover $u\mapsto t^{2}$. In a neighborhood of $\infty$, we can express the projection map $X\times_{\mathbb{P}^{1}_{u}}\mathbb{P}^{1}_{t}\to X$ as
@@ 3.1.6. Chains functor
[…]
@@ 3.3. Localization and completion
$\operatorname{Mod}_{\operatorname{Sp}}(R[r^{-1}])$ is defined as the presentable localization (or Bousfield localization) of the $\infty$-category $\operatorname{Mod}_{\operatorname{Sp}}R$ away from $R/r=\operatorname{cof}(\Sigma^{i}R\xrightarrow{r}R)$. In other words, it is the full subcategory of $\operatorname{Mod}_{\operatorname{Sp}}(R)$ such that $\operatorname{map}(R/r,-)$ vanishes, and the left adjoint of the inclusion of this subcategory sends $M$ to $M[r^{-1}]$.
@@
There are natural $\mathbb{E}_{1}$-ring maps $\to R^{\wedge}_{r}, RR$$\to R[r^{-1}$] induced by taking the map induced on endomorphism rings of the $R$-module $R$ via the completion and localization functor. Note also that $R[r^{-1}]=R[(r^{n})^{-1}]$ and $R^{\wedge}_{r}=R^{\wedge}_{r^{n}}$ for each $n\geq 1$.
@@
Because $x^{m}$ maps to 0 in the component $R[(1-e)^{-1}]\simeq R^{\wedge}_{x}$, the kernel of surjective map $\pi_{*}R\to\pi_{*}R^{\wedge}_{x}$ contains two sided ideal generated by $x^{m}$. From the above, this kernel is equal to $\pi_{*}R[x^{-1}]$. Since $\pi_{*}R[x^{-1}]$ is contained in the two sided ideal generated by $x^{m}$ since any $y\in\pi_{*}R[x^{-1}]$ can be written as $yx^{-m}x^{m}$, we obtain that the kernel is equal to the two sided ideal generated by $x^{m}$. ∎
@@ 3.4. Tools for homological epimorphisms
Using § 3.4, we learn that $S\to S^{\prime}$ is a homological epimorphism. Letting $I$ be the fiber of $S\to S^{\prime}$, we have $\otimes_{S}S^{\prime}=0. II$ is the fiber of a map of connective $\mathbb{E}_{1}$-rings, so it is 1-1-connective. We will inductively show on $i\geq-1$ that $\pi_{i}I$ is 0. The inductive hypothesis implies that $I$ is $i$-connective, so and since $S\to S^{\prime}$ is an equivalence on $\pi_{0}$, we claim
@@ 3.5. Nilpotence of actions
${X^{\prime}}X′′{X^{\prime\prime}}$${Y^{\prime}}Y′′.{Y^{\prime\prime}.}$$\scriptstyle{f}$$\scriptstyle{0}$$\scriptstyle{f^{\prime}}$$\scriptstyle{f^{\prime\prime}}$$\scriptstyle{0}$
@@
In particular, if a natural endomorphism $g$ of the identity functor on $\mathcal{C}$} acts by 0 on $X$ and $Y$, then $g^{2}$ acts by 0 on $\operatorname{cof}(f)$ for any map $X\to Y$.
@@
For the first claim, the composite $\operatorname{cof}(f)\to\operatorname{cof}(f^{\prime})\to\Sigma X^{\prime}$ is null since $X\to X^{\prime}$ is null, so we obtain a factorization of $\operatorname{cof}(f)\to\operatorname{cof}(f^{\prime})$ through $Y^{\prime}$. Composing with the second horizontal map, since $Y^{\prime}\to Y^{\prime\prime}$ is 0, we obtain a nullhomotopy of the desired map.
For the second claim, we apply the first claim when $X=X^{\prime}=X^{\prime\prime},Y=Y^{\prime}=Y^{\prime\prime}$ and $f=f^{\prime}=f^{\prime\prime}$ and the squares come from the natural endomorphism $g$. ∎
@@
Let $\mathcal{C}$} be an exactly monoidal stable $\infty$-category, let $f:\operatorname{\mathbb{1}}\to\Sigma^{i}\operatorname{\mathbb{1}}$ be a map in $\mathcal{C}$}, where $\operatorname{\mathbb{1}}$ is the monoidal unit. Then the map $\operatorname{cof}(f)\otimes f^{\otimes 2}$ is null.
In particular, if $R$ is an $\mathbb{E}_{1}$-algebra, taking $\mathcal{C}$} to be the category of $R$-bimodules, if $f:R\to\Sigma^{i}R$ a bimodule map, and $M\in\operatorname{Mod}_{\operatorname{Sp}}(R)$, then $f^{\otimes 2}$ acts by 0 on $\operatorname{cof}f\otimes_{R}M$ as a map of $R$-modules.
@@
We turn to proving the first statement. Let $\delta:\operatorname{cof}(f)\to\Sigma^{1}\operatorname{\mathbb{1}}$ be the boundary map in the cofiber sequence. The map $\delta\circ(\operatorname{cof}(f)\otimes f)$ is null since it agrees with $\operatorname{cof}(f)\xrightarrow{\delta}\Sigma^{1}\operatorname{\mathbb{1}}\xrightarrow{f}\Sigma^{1+i}\operatorname{\mathbb{1}}$. It follows that $\operatorname{cof}(f)\otimes f$ factors through $\operatorname{\mathbb{1}}$. But then since $\operatorname{\mathbb{1}}\xrightarrow{f}\Sigma^{i}\operatorname{\mathbb{1}}\to\operatorname{cof}(f)$ agrees with $\operatorname{\mathbb{1}}\to\Sigma^{-i}\operatorname{cof}(f)\xrightarrow{f}\operatorname{cof}(f)$, it follows that $f^{\otimes 2}$ vanishes on $\operatorname{cof}(f)$. ∎
@@ 4.1. Background on racks
We can identify the connected components of the pointed Hurwitz space $\operatorname{Hur}^{c}_{n}$ with orbits of the $B_{n}$ action on $c^{n}$. Under this identification, for $x_{1},\ldots,x_{n}\in c$ we use $x_{1}]\cdots[x_{n}]$ to denote the connected component of $\operatorname{Hur}^{G,c}_{n,\mathbb{C}}$ corresponding to the $B_{n}$ orbit of the tuple $(x_{1},\ldots,x_{n})$. We will also use $\alpha_{x}$ as equivalent notation for $x]\in\pi_{0}\operatorname{Hur}^{G,c}_{n,\mathbb{C}}$. See also § 4.4.
@@ 4.2. The main stability result
First, note that there is an identification between the pointed Hurwitz space $\operatorname{Hur}_{n}^{G,c}$ and $\operatorname{Hur}_{n}^{c}$ as both can be described as the homotopy quotient $(c^{n})_{hB_{n}}$, see § 1.4.2. Let us fix a value of $i\geq 0$. By [EVW16, Theorem 6.1], there exist constants $I,J$, and DD [^4] In fact, one may take $D=1$ by [EL23, Proposition A.3.1]. so that for any $i\geq 0$ and $n>Ii+J$ and $U:=\sum_{h\in c}\alpha_{h}^{D\operatorname{ord}(h)}$, $\pi_{i}A_{n}$ agrees with $\pi_{i}A[U^{-1}]_{n}$.
@@ 4.3. Centrality
Informally, in the case of Hurwitz spaces associated to groups, the reason for this stronger form of centrality is that there is a homotopy $\alpha][\beta]\simeq[\beta][\beta^{-1}\alpha\beta]$ obtained by changing the direction from which one concatenates $G$-covers. Thus if $\beta$ centralizes $c$, this identifies left and right multiplication by $\beta$. The above argument is not enough to make multiplication by $\beta$ into a bimodule map, but it can be made as such. We now argue this below and further generalize from groups to racks:
@@
Let $m$ be such that $\gamma\in\operatorname{Hur}_{m}^{c}$. Consider $m_{!}(*)$, where $m$ is considered as the inclusion of the base point $\to BB_{m}$. Explicitly, $m_{!}(*)$ is the local system which is $B_{m}$ in degree $m$ and empty otherwise. There is a left module map $(c^{i})_{i\in\mathbb{N}}\otimes m_{!}(*)\to(c^{i})_{i\in\mathbb{N}}$, given by right multiplying by the class $\gamma$ on the identity element of $B_{m}$. Here we realize $(c^{i})_{i\in\mathbb{N}}\otimes m_{!}(*)$ as a bimodule via the $\mathbb{E}_{2}$-monoidal structure, i.e using the twist map $(c^{i})_{i\in\mathbb{N}}\otimes m_{!}(*)\to m_{!}(*)\otimes(c^{i})_{i\in\mathbb{N}}$. The fact that this multiplication map is a bimodule map follows if the diagram
@@
Let $\gamma\in\pi_{0}\operatorname{Hur}^{c}$ be central. Then the associated class $\gamma]\in\pi_{0}A$ is $\mathbb{E}_{2}$-central.
@@ 4.4. Decomposition of the comparison map
Let $c$ be a finite rack. Given an element $h\in c$, let us use $\alpha_{h}$ as alternate notation for the element $[h]$ in $\pi_{0}\operatorname{Hur}_{1}^{c}$, as defined in § 4.1. For a subset $S\subset c$, we use $\alpha_{S}$ to denote the set $\{\alpha_{s}:s\in S\}$. We use $M[\alpha_{S}^{-1}]$ to denote the localization of some $A$-module $M$ at the set $\{\alpha_{s},s\in S\}$ and we similarly use $M^{\wedge}_{\alpha_{S}}$ to denote the completion of $M$ at $\{\alpha_{s}:s\in S\}$ (when this doesn’t depend on an order of $S$).
@@
With notation as in § 4.4, let $c$ be a finite nonempty non-splitting rack and let $c^{\prime}\subset c$ be a subrack. In $A[\alpha_{c^{\prime}}^{-1}]$, any element $\alpha_{\beta}$ with $\beta\in c-c^{\prime}$ satisfies the conditions of § 3.3. Thus for an AA-module $M$, we can form $M[\alpha_{c^{\prime}}^{-1}]^{\wedge}_{\alpha_{c-c^{\prime}}}$ as in § 3.3.
@@
is an equivalence, where $\coprod_{i=1}^{|c-c^{\prime}|}\beta_{i}=c-c^{\prime}$. Here, we have used § 3.4 to commute the localization with the quotienting, and § 4.3 to make sense of the iterated quotient. Namely, because $\alpha_{\beta_{i}}^{\operatorname{ord}(\beta_{i})}$ is a bimodule map by § 4.3, we can tensor with the AAA-A-bimodule $(A/\alpha^{\operatorname{ord}(\beta_{1})}_{\beta_{1}})$ to make sense of the tensor product above as an $A$-module.
Note that on the source and target of this map, for each $\beta\in c-c^{\prime}$, $\alpha_{\beta}^{\operatorname{ord}(\beta)n_{\beta}}$ acts by 0 for some $n_{\beta}$. This follows from applying § 3.5 to see that after quotienting out by $\alpha_{\beta}^{\operatorname{ord}(\beta)}$, $\alpha_{\beta}^{2\operatorname{ord}(\beta)}$ acts by 0, and using § 3.5 to see that after each further quotient, the smallest power of $\alpha_{\beta}^{\operatorname{ord}(\beta)}$ that acts by 0 gets multiplied by at most $2$.
If $c^{\prime}$ is empty, then we claim that both the source and target of the map (4.1) are 0. We claim that some power of $U$ acts by 0 on the source and target, since a large enough power of $U$, $U^{j}$, is in the two sided ideal of $\pi_{0}A$ generated by $\alpha_{\beta}^{\operatorname{ord}(\beta)n_{\beta}}$. This is because by the pigeonhole principle, one can see that any element of $\pi_{0}\operatorname{Hur}^{G,c}_{n}$, written as a product of elements in $c$ for $n$ large enough contains $\max_{\beta\in c}(\operatorname{ord}(\beta)n_{\beta})$ many copies of a single $\beta\in c$. By pulling these to the left using the braid group action, we see the element is divisible by $\alpha_{\beta}^{\operatorname{ord}_{\beta}n_{\beta}}$. Since $U^{j}$ is in the ideal generated by the smallest powers of $\alpha_{\beta}^{\operatorname{ord}(\beta)}$ that act trivially, and hence it acts trivially. Since $U$ acts both as an isomorphism and nilpotently, the source and target of (4.1) with $c^{\prime}=\emptyset$ are both 0.